### v1

# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

Abstract
This paper focuses on the problem of automatic link selection in multi-channel multiple access control using bandit feedback. In particular, a controller assigns multiple users to multiple channels in a time slotted system, where in each time slot at most one user can be assigned to a given channel and at most one channel can be assigned to a given user. Given that user $i$ is assigned to channel $j$, the transmission fails with a fixed probability $f_{i,j}$. The failure probabilities are not known to the controller. The assignments are made dynamically using success/failure feedback. The goal is to maximize the time average utility, where we consider an arbitrary (possibly nonsmooth) concave and entrywise nondecreasing utility function. The problem of merely maximizing the total throughput has a solution of always assigning the same user-channel pairs and can be unfair to certain users, particularly when the number of channels is less than the number of users. Instead, our scheme allows various types of fairness, such as proportional fairness, maximizing the minimum, or combinations of these by defining the appropriate utility function. We propose two algorithms for this task. The first algorithm is adaptive and gets within $\mathcal{O}(\log(T)/T^{1/3})$ of optimality over any interval of $T$ consecutive slots over which the success probabilities do not change. The second algorithm has faster $\mathcal{O}(\sqrt{\log(T)/T})$ performance over the first $T$ slots, but does not adapt well if probabilities change.
Index Terms:
Multi-armed bandit learning; Proportional fairness; Network utility maximization; Optimization; Stochastic control; Adaptive learning
I Introduction
We consider the Multiple Access Control (MAC) problem with $n$ users and $m$ channels in slotted time $t\in\mathbb{N}$. In each time slot, a controller has to assign the users to channels such that at most one user is assigned to given channel and at most one channel is assigned to a given user. The channel assignments may fail. In particular, there exist $q_{i,j}\in[0,1]$ for each $(i,j)\in\{1,2,\dots,n\}\times\{1,2,\dots,m\}$, where in time slot $t$, given that the controller decided to assign user $i$ to channel $j$, the assignment fails independently with probability $1-q_{i,j}$. The controller does not know the probabilities $q_{i,j}$. Instead, at the end of every slot, it receives feedback on whether the transmission for each assigned user-channel pair succeeded or failed.
Define the matrices $\boldsymbol{Y}(t),\boldsymbol{S}(t)\in\{0,1\}^{n\times m}$ and vector $\boldsymbol{X}(t)\in\{0,1\}^{n}$, where
$S_{i,j}(t)=\begin{cases}1&\text{ if link }i,j\text{ is successful in time slot }t\\ 0&\text{ otherwise},\end{cases}$
$Y_{i,j}(t)=\begin{cases}1&\text{ user }i\text{ is assigned to channel }j\text{ in time slot }t\\ 0&\text{ otherwise},\end{cases}$
and $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for all $i\in\{1,2,\dots,n\}$. The goal is to maximize $\lim_{T\to\infty}\phi(\mathbb{E}\{\boldsymbol{\overline{X}}(T)\})$ using feedback on the link failures, where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a concave entrywise nondecreasing utility function known to the controller and $\boldsymbol{\overline{X}}(T)=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)$.  [^1] The limit is assumed to exist for simplicity of this introduction; the precise goal is to maximize a $\lim\inf_{T\to\infty}\phi(\boldsymbol{\overline{X}}(T))$
We also focus on establishing finite time bounds. In particular, for given a finite time horizon $T\in\mathbb{N}$, we require the algorithm to satisfy
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)\leq g(T),$
where $\phi^{\text{opt}}$ is the optimal utility of the original problem and $g$ is a nonnegative function such that
$\lim_{T\to\infty}g(T)=0$
In addition, we are looking for algorithms that are adaptive. Formally, consider a system in which the channel success probabilities may change. In such a system, given a $T\in\mathbb{N}$, we require
$\phi^{\text{opt}}-\phi\left(\frac{1}{T}\mathbb{E}\left\{\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)\leq g(T)$
for any $T_{0}\in\mathbb{N}$, irrespective of the success probabilities outside of the time frame $[T_{0}:T_{0}+T-1]$, given that success probabilities remained constant in the frame. Here, $\phi^{\text{opt}}$ is the optimal utility of the original problem that uses the constant success probabilities in $[T_{0}:T_{0}+T-1]$ of the above scenario. Note that $g$ is the same function regardless of $T_{0}$.
I-A Utility functions
One utility function that can be used in our work is $\phi(\boldsymbol{x})=\min\{x_{1},\dots,x_{n}\}$. This is a nonsmooth utility function that seeks to maximize the minimum time-average success rate across all users. However, if there is one user with very low success probability, this utility function can cause almost all the resources to be devoted to that user, resulting in poor performance for all users.
Another choice is $\phi(\boldsymbol{x})=w_{1}\min\{x_{1},\dots,x_{n}\}+w_{2}\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $w_{1},w_{2},\beta$ are given nonnegative weights. The logarithmic term introduces a form of proportional fairness [1, 2]. See also discussion of different utility functions in [1, 2, 3, 4, 5]
I-B Related work
Network scheduling in stochastic environments is a widely considered problem in the literature on communication networks. Some examples include scheduling in wireless networks [6, 7], computer networks [8], vehicular networks [9] and unmanned aerial vehicle networks [10]. These works consider different goals such as optimal scheduling for queue-stability [11], power minimization [12], utility maximization [13], ensuring fairness [14], flow control [15], and multiple access control [16].
Multiple access, where many users access a limited number of communication channels, is an important problem in network scheduling. Here, it is desirable for users to be scheduled to avoid collisions. Link failures occur when the receiver is unable to decode the packet transmissions. This can occur, for example, when a fixed transmission power is used, but channel conditions have random and unknown fluctuations so that the received signal strength is insufficient for decoding. Different links can have different properties (such as different geographic distances to the receiver), so they can also have different success probabilities. The naive approach of assigning the users with links with the least failure probabilities leads to unfairness since, in such a scenario, users with high link failure probabilities will never be assigned. A common approach to solve this problem is to maximize a utility function of the time-averaged success. This problem has been well studied in the full information scenario when either the fluctuating channel conditions are known before transmission (called opportunistic scheduling) or when channel success probabilities are known in advance. Opportunistic scheduling has been considered using utility functions [17], Lyapunov drift [18], Frank-Wolfe [19, 20, 21], primal-dual [22, 23], and drift-plus-penalty [24]. The case when success probabilities are known in advance can be solved offline as a convex optimization problem using the mirror descent technique [25].
The problem becomes challenging when the success probabilities are not known, but we only receive bandit feedback on the successes. The problem has to be approached by combining ideas from optimizing functions of time averages with multi-armed bandit learning. The work on bandits with vector rewards and concave utility functions can be adapted for the single channel case ($m=1$) of our problem (See [26, 27, 28]). There are two main drawbacks to these approaches. First, the above works do not consider the matching constraints considered in our work. Next, they focus on upper confidence bound (UCB) techniques and are not adaptive. It is possible to develop adaptive algorithms using techniques from the above works together with the EXP3 algorithm [29] for the single channel case. However, they cannot be directly extended for the multi-channel case. The main reason for this is the complexity of the inner problem arising in each iteration which is a problem over the set of doubly stochastic matrices (Birkhoff polytope). This can be addressed using Sinkhorn’s algorithm [30] in each iteration. However, this results in an algorithm that requires computationally complex inner iterations to be performed within each iteration. The work of [31] uses follow the regularized leader approach to solve adversarial bandit problems over the set of doubly stochastic matrices. However, their algorithm relies on a computationally expensive inner iteration similar to what we described before. The work of  [32] proposes an algorithm to solve online optimization problems over transport polytopes. All the inner iterations of their algorithm have explicit solutions thanks to the rounding trick introduced for transport polytopes by [33]. In our work, we adapt the rounding trick to the Birkhoff polytope and develop an algorithm for general $m$. We also develop a UCB-based algorithm for our problem. It should be noted that although this algorithm achieves faster convergence, it is not adaptive. However, the UCB algorithm has additional advantages such as distributed implementation. Our paper also treats general (possibly nonsmooth) utility functions, so Frank-Wolfe methods that rely on smoothness cannot be used.
Existing work on using learning for scheduling problems in communications include spectrum sharing [34, 35], network utility maximization with unknown utility functions [36], and queue stability [11]. The existing literature on bandit learning-based methods for scheduling problems focuses on maximizing throughput [37]. The work of [36] considers bandit learning with unknown utilities, and the work of [38] considers the online resource allocation problem, where the objective is to maximize the $\alpha$-fairness of the allocations. Neither of the above works considers the bandit reward structure, and both works have a different problem structure from ours. In addition, the above works do not focus on the adaptiveness of the proposed algorithms.
It should be noted that the concept of adaptiveness considered in this paper is different from the adversarial setting [39, 40, 29], although they have similarities. In adversarial settings, optimality is defined with respect to the rewards received throughout the time horizon. This is useful when no stochastic assumptions can be placed on the rewards. On the other hand, the adaptiveness considered in this paper is useful when we can place stochastic assumptions on the rewards, but the distributions may change from time to time. In this case, we can define an optimal strategy for each time frame in which the reward distribution remained constant. Hence, the goal in this frame is to learn the aforementioned strategy irrespective of the reward distributions of the past. An approach commonly used in problems of this flavor is minimizing dynamic regret [41, 42, 43]. Here, the regret is modified to account for the changing environments and the regret bounds are in terms of some measure that captures the degree of change. Various algorithms are developed for the setting with linear utility functions using optimizing in phases/episodes [41, 42], and sliding window-based algorithms [43]. We utilize a simpler notion of adaptiveness and develop an algorithm for the case with general utility functions.
Multi-armed bandit (MAB) learning [44, 45] is a class of extensively studied problems. In the classical MAB setting, a user chooses from multiple arms, each incurring a fixed mean cost. The choice is made without knowledge of the mean costs, using the feedback received on the cost of the chosen arm. The goal is to learn the arm with the lowest mean cost as fast as possible. Typically, the algorithms implemented to solve such problems involve an exploration phase, where the user explores all arms in order to learn the mean costs, after which, in the exploitation phase, the user exploits the learned information to make good decisions [46, 47]. Upper confidence bound-based algorithms, where the algorithm maintains an upper bound on the mean cost of each arm, is a popular approach to solve these problems [47, 48]. Adversarial bandit learning [39, 40], where an adversary is allowed to assign the cost of each arm after the user selects the arm, is an extension of multi-armed bandit learning. In such problems, no stochastic assumptions are placed on the rewards. These problems are solved using algorithms such as the EXP3 algorithm [29], where in each iteration of the algorithm, a probability distribution over the space of all possible actions is found, after which the decision is randomly sampled from the distribution. There are also many extensions of the MAB problem, such as linear bandits, contextual bandits, and combinatorial bandits [47].
I-C Our Contributions
We develop and analyze two algorithms to solve the problem of automatic link selection in multi-channel multiple access, combining the ideas of multi-armed bandit learning and Lyapunov optimization. Although the classical MAB problem has been widely analyzed for maximizing linear utilities, they suffer from lack of fairness in assignments when applied to the considered problem. It is notable that our method allows either smooth or nonsmooth concave, entrywise nondecreasing utilities.
We prove that the first algorithm gets within $\mathcal{O}(T^{-1/3}\log(T))$ of optimality over any interval of $T$ consecutive time slots during which the (unknown) success probabilities do not change. If these probabilities are different before $T_{0}$, but change to new probabilities during $\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, our performance guarantees for the new interval are independent of behaviors before $T_{0}$, even though the algorithm does not know the exact time $T_{0}$ of the change. Hence, the algorithm is adaptive. This is possible thanks to the importance sampling technique in [29], which we use to estimate the true link failure probabilities in each time slot. This technique uses feedback only from the previous time slot.
The second algorithm is based on upper confidence bound (UCB) techniques and gets within $\mathcal{O}(\sqrt{\log(T)/T})$ of optimality within a finite time horizon of $T$ time slots. Although this algorithm has faster convergence compared to the first algorithm, it is not adaptive due to the UCB-based estimation. However, the channel-user assignment of this algorithm has a max-weight structure and can be solved using the well-known Hungarian algorithm [49]. This is in contrast to the first algorithm, where the assignment is sampled from a distribution in each time slot, which may be less desirable in certain scenarios. Due to the max-weight decisions, the second algorithm can be implemented in a distributed setting in the absence of a centralized controller, given that all the users have access to feedback on which channels are successfully accessed in each time slot. Simulations also depict the adaptiveness of the first algorithm and the faster convergence of the second algorithm.
I-D Notation
[…]
I-E Definitions
In this subsection, we define some quantities that will be useful throughout. Define $s=\max\{n,m\}$. Also, define the sets
$\Delta_{l,\varepsilon}=\Big{\{}\boldsymbol{p}\in\mathbb{R}^{k}:\sum_{i=1}^{l}p_{i}=1,p_{i}\geq\varepsilon\ \forall i\in[l]\Big{\}},\text{ where }l\in\mathbb{N},$
$\mathcal{S}^{\text{row}}_{\varepsilon}=\Bigg{\{}\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}:\sum_{k=1}^{s}P_{i,k}=1,P_{i,j}\geq\varepsilon\ \forall i,j\in[s]\Bigg{\}},$
$\mathcal{S}^{\text{col}}_{\varepsilon}=\Bigg{\{}\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}:\sum_{k=1}^{s}P_{k,j}=1,P_{i,j}\geq\varepsilon\ \forall i,j\in[s]\Bigg{\}},$
$\mathcal{S}^{\text{doub}}_{\varepsilon}=\mathcal{S}^{\text{col}}_{\varepsilon}\cap\mathcal{S}^{\text{row}}_{\varepsilon},\ \mathcal{S}_{\varepsilon}=\mathcal{S}^{\text{col}}_{\varepsilon}\cup\mathcal{S}^{\text{row}}_{\varepsilon}.$
We also denote $\Delta_{l}=\Delta_{l,0}$, $\mathcal{S}^{\text{row}}=\mathcal{S}^{\text{row}}_{0}$, $\mathcal{S}^{\text{col}}=\mathcal{S}^{\text{col}}_{0}$, $\mathcal{S}^{\text{doub}}=\mathcal{S}^{\text{doub}}_{0}$, and $\mathcal{S}=\mathcal{S}_{0}$. We hide the dependence on $s$ in the notation for sets for clarity.
I-F Assumptions
Before moving on to the problem, we state our main assumptions.
  • A1
    The function $\phi$ is concave, entrywise nondecreasing, and has bounded subgradients in $[0,1]^{n}$, i.e, $|\phi^{{}^{\prime}}_{i}(\boldsymbol{x})|\leq B\ \forall i\in[n]$, and $\boldsymbol{x}\in[0,1]^{n}$. Hence, $\phi$ is $\sqrt{n}B$-Lipschitz continuous. Also, let $\phi^{\text{max}}=\max_{\boldsymbol{x}\in[0,1]^{n}}|\phi(\boldsymbol{x})|$.
  • A2
    We have access to the solution of the problem $\max_{\boldsymbol{x}\in[0,1]^{n}}[\phi(\boldsymbol{x})+\sum_{i=1}^{n}c_{i}x_{i}]$, for all $\boldsymbol{c}\in\mathbb{R}^{n}_{+}$.
    Note: This assumption is valid for most separable functions $\phi$. For instance, when $\phi$ is a proportionally fair utility function type of the form $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta\in\mathbb{R}_{+}$, the problem has an explicit solution.
II Problem Setup
We formalize the problem of interest below.
(P1:) $\underset{\boldsymbol{Y}(1),\boldsymbol{Y}(2),\dots}{\text{max}}\lim\inf_{T\to\infty}\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$ (1)
[…]
$\ \ \ \ \ \ \ t,\tau\in\mathbb{N}\text{ and }\tau\geq t$ (2)
$\ \ \ \ \ \boldsymbol{Y}(t),\text{ and }S_{i,j}(\tau)\text{ are independent for all }t,\tau\in\mathbb{N},$
$\ \ \ \ \ \ \ (i,j)\in[n]\times[m],\tau<t,Y_{i,j}(\tau)\neq 1$ (3)
$\ \ \ \ \ \boldsymbol{Y}(t)\in\{0,1\}^{n\times m}\ \forall t\in\mathbb{N}$ (4)
$\ \ \ \ \ \sum_{i=1}^{n}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},j\in[m]$ (5)
$\ \ \ \ \ \sum_{j=1}^{m}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},i\in[n]$ (6)
$\ \ \ \ \ X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)\forall t\in\mathbb{N},\ i\in[n],$ (7)
where constraint (2) ensures transmission decisions do not know success/failures before they happen; (3) ensures we cannot use information that is never observed. Define $s=\max\{n,m\}$ and $\phi^{\text{opt}}$ as the optimal objective value of (P1).
Lemma 1
Consider the following problem.
(P2:) $\underset{\boldsymbol{P},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (8)
[…]
$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (10)
$\ \ \ \ \ \sum_{j=1}^{m}q_{i,j}P_{i,j}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (11)
and let $\phi^{*}$ denote the optimal objective value of the above problem. Then we have $\phi^{*}=\phi^{\text{opt}}$.
Proof:
We first prove that $\phi^{*}\geq\phi^{\text{opt}}$. Fix $\varepsilon>0$. Then there is a positive integer $T$ and decisions $\boldsymbol{Y}^{*}(1),\boldsymbol{Y}^{*}(2),\dots,\boldsymbol{Y}^{*}(T)$ that respect constraints (2)-(6), such that
$\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}^{*}(t)\}\right)\geq\phi^{\text{opt}}-\varepsilon,$ (12)
where $X^{*}_{i}(t)=\sum_{j=1}^{m}Y^{*}_{i,j}(t)S_{i,j}(t)$ for all $t\in\mathbb{N}$ and $i\in[n]$. Define the matrix $\boldsymbol{\tilde{P}}^{*}\in\mathbb{R}^{s\times s}$ such that
$\tilde{P}^{*}_{i,j}=\begin{cases}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y^{*}_{i,j}(t)\}&\text{ if }(i,j)\in[n]\times[m]\\ 0&\text{ otherwise. }\end{cases}$
Let us define $\boldsymbol{P}^{*}\in\mathcal{S}^{\text{doub}}$ such that $\boldsymbol{P}^{*}\geq\boldsymbol{\tilde{P}}^{*}$. It can be observed that such a $\boldsymbol{P}^{*}$ exists from the definition of $\boldsymbol{\tilde{P}}^{*}$ and the constraints (4)-(6). Let $\boldsymbol{\gamma}^{*}\in[0,1]^{n}$ such that $\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}$ for $i\in[n]$. It is easy see that $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ satisfy constraints (9)-(11). Also, we have
$\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}\geq\sum_{j=1}^{m}q_{i,j}\tilde{P}^{*}_{i,j}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{X^{*}_{i}(t)\}$ (13)
since (2) ensures $\mathbb{E}\{Y^{*}_{i,j}(t)S_{i,j}(t)\}=\mathbb{E}\{Y^{*}_{i,j}(t)\}q_{i,j}$. Hence,
$\phi^{*}$ $\geq_{(a)}\phi(\boldsymbol{\gamma}^{*})\geq_{(b)}\phi\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{\boldsymbol{X}^{*}(t)\}\right)\geq_{(c)}\phi^{\text{opt}}-\varepsilon,$
where (a) follows since $\phi^{*}$ is the optimal objective value of (P2) and $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is feasible for (P2), (b) follows from (13) and the entrywise nondecreasing property of $\phi$ (see assumption A1) and (c) follows from (12). The above is true for all $\varepsilon>0$. Hence, we have $\phi^{*}\geq\phi^{\text{opt}}$ as desired.
Now, we prove that $\phi^{*}\leq\phi^{\text{opt}}$. Let $\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*}$ denote the optimal solution for (P2). Using Birkhoff-von Neumann Theorem [50], notice that there exists $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}^{*}=\sum_{l=1}^{r}s_{l}\boldsymbol{M}^{l}$, and $\boldsymbol{s}\in\Delta_{r}$. In each time slot, we simply sample $l_{t}\sim\boldsymbol{s}(t)$ and set $\boldsymbol{Y}(t)$ according to $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for $i\in[n],j\in[m]$. It can be easily seen that this policy yields an objective value of $\phi^{*}$ for (P1). Hence, we are done. ∎
III Adaptive Algorithm
Now, we develop our first algorithm. The idea is to find $\boldsymbol{\tilde{P}}(t)$ as a column stochastic matrix and a row stochastic matrix alternatively in odd and even slots. Then we obtain $\boldsymbol{P}(t)$ by approximating $\boldsymbol{\tilde{P}}(t)$ by a doubly stochastic matrix, using the rounding trick similar to the one introduced in [33], after which we sample a permutation matrix from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [50]. To do the assignment in time slot $t-1$, we discard the last $s-n$ rows or last $s-m$ columns of the permutation matrix, depending on whether $m$ or $n$ is larger.
1 Initialize $\boldsymbol{\tilde{P}}(1)\in\mathcal{S}^{\text{doub}}_{\varepsilon}$ and the virtual queues $\boldsymbol{Q}(1)\in[0,BV+1]^{n}$ arbitrarily.
2 for each time slot $t\in\mathbb{N}$ do
3       Set $\boldsymbol{P}(t)=\text{ROUND}(\boldsymbol{\tilde{P}}(t))$ (This function yields $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ ).
4
5      Using Birkhoff-von Neumann Decomposition [50], find $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}(t)=\sum_{l=1}^{r}s_{l}(t)\boldsymbol{M}^{l}$, and $\boldsymbol{s}(t)\in\Delta_{r}$.
6       Sample $l_{t}\sim\boldsymbol{s}(t)$ and take action $\boldsymbol{Y}(t)$, where $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for all $i\in[n]$, $j\in[m]$, and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
7       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i,j}(t)=S_{i,j}(t)Y_{i,j}(t)/P_{i,j}(t)$ for all $i\in[n]$, and $j\in[m]$.
8       Find $\boldsymbol{\gamma}(t+1)\in[0,1]^{n}$, and $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{Q}$ using
$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big{]}.$
$\boldsymbol{\tilde{P}}(t+1)=\arg\min_{\boldsymbol{P}\in\mathcal{Q}}\Big{[}-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\Big{]},$ (14)
where $\mathcal{Q}=\mathcal{S}^{\text{col}}_{\varepsilon}$ if $t$ is even, and $\mathcal{Q}=\mathcal{S}^{\text{row}}_{\varepsilon}$ if $t$ is odd, and the divergence $D$ is defined by
$D(\boldsymbol{X}\lVert\boldsymbol{Y})=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right).$ (15)
for all $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}.$
9       Update the virtual queues
$\boldsymbol{Q}(t+1)=\left[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right]_{+},$ (16)
where $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$ for $i\in[n]$.
10
Algorithm 1 Adaptive MAC
Below, we introduce the ROUND function, a technique adapted from the one introduced in [33] to approximate a nonnegative matrix by a matrix in the transport polytope. The technique can be readily extended to the Birkhoff polytope using the same algorithm.
ROUND($\boldsymbol{P}$) function for $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$ :
  1. 1.
    Define the matrix $\boldsymbol{P}^{{}^{\prime}}$ (row normalization of $\boldsymbol{P}$) using
    $P^{{}^{\prime}}_{i,j}=\begin{cases}\frac{P_{i,j}}{\sum_{l=1}^{s}P_{i,l}}&\text{ if }\sum_{l=1}^{s}P_{i,l}>1\\ P_{i,j}&\text{ otherwise.}\end{cases}$ (17)
  2. 2.
    Define the matrix $\boldsymbol{P}^{{}^{\prime\prime}}$ (column normalization of $\boldsymbol{P}^{{}^{\prime}}$) using
    $P^{{}^{\prime\prime}}_{i,j}=\begin{cases}\frac{P^{{}^{\prime}}_{k,j}}{\sum_{k=1}^{s}P^{{}^{\prime}}_{k,j}}&\text{ if }\sum_{k=1}^{s}P^{{}^{\prime}}_{k,j}>1\\ P^{{}^{\prime}}_{i,j}&\text{ otherwise.}\end{cases}$ (18)
  3. 3.
    Define the output matrix $\boldsymbol{Q}$,
    $\boldsymbol{Q}=\begin{cases}\boldsymbol{P}^{{}^{\prime\prime}}+\frac{(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}}{C}&\text{ if }C\neq 0\\ \boldsymbol{P}^{{}^{\prime\prime}}&\text{ otherwise, }\end{cases}$ (19)
    where $C=\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}$.
It can be shown that ROUND$(\boldsymbol{P})\in\mathcal{S}_{\varepsilon/s}^{\text{doub}}$ whenever $\boldsymbol{P}\in\mathcal{S}_{\varepsilon}$.
In Algorithm 1, we provide the algorithm for the task. In the following section, we focus on solving the intermediate problem (8), after which we move on to the analysis of the Algorithm. Problem (P2) motivates the definition of auxiliary variables $\boldsymbol{\gamma}(1),\boldsymbol{\gamma}(2),\dots\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(1),\boldsymbol{Q}(2),\dots$ to aid the algorithm. In particular, we will show that we can use our algorithm to get arbitrarily close to $\phi^{\text{opt}}$ in objective value.
III-A Solving Problem (8)
Finding $\boldsymbol{\gamma}(t+1)$: Notice that this problem can be solved due to the Assumption A2.
Finding $\boldsymbol{\tilde{P}}(t+1)$: We will only consider the case when $t$ is even. The case when $t$ is odd can be solved similarly. Notice that we can separately solve for each column of $\boldsymbol{\tilde{P}}(t+1)$. To solve for the $j$-th column of $\boldsymbol{\tilde{P}}(t+1)$, we define $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{s}$, where
$x_{i}=\begin{cases}\eta Q_{i}(t)\hat{S}_{i,j}(t)&\text{ if }i\in[n],j\in[m]\\ 0&\text{ otherwise}\end{cases}$
and $\boldsymbol{y}$ is the $j$-th column of $\boldsymbol{\tilde{P}}(t)$. The problem to be solved is
(P3:) $\underset{\boldsymbol{p}}{\text{min}}-\sum_{i=1}^{s}x_{i}p_{i}+D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{y})$
[…]
where $D_{\text{KL}}(\cdot\lVert\cdot)$ in this case is the KL-divergence. It should be noted that (P3) has a classic structure that is solved in [11]. In particular, we define $\boldsymbol{z}$, where $z_{i}=y_{i}\exp(x_{i})$. First, assume that $\boldsymbol{z}$ is sorted in the increasing order. Then it can be shown that there exists $i\in[0:s-1]$ such that the vector $\boldsymbol{u}^{i}\in\mathbb{R}^{s}$ given by,
$u^{i}_{j}=\begin{cases}\varepsilon&\text{ if }j\leq i\\ \frac{z_{j}}{\sum_{l=i+1}^{s}z_{l}}(1-\varepsilon i)&\text{ if }j>i\\ \end{cases}$
satisfies, $u^{i}_{j}\geq\varepsilon$ for all $j\in[i+1:s]$. Then it can be shown that $\boldsymbol{u}^{i}$ is the solution to (P3). Hence, solving (P3) amounts to calculating $\boldsymbol{u}^{i}$ for each $i\in[0:s-1]$ and checking the above condition.
We first establish the following two lemmas regarding the ROUND operation. A version of the following lemma is proved for the transport polytope in [33]. We adapt the idea for the Birkhoff polytope.
Lemma 2
The input $\boldsymbol{P}\in\mathbb{R}^{s\times s}_{+}$, intermediate matrices $\boldsymbol{P}^{{}^{\prime}},\boldsymbol{P}^{{}^{\prime\prime}}$ and the output matrix $\boldsymbol{Q}$ of the ROUND function in Algorithm 1 satisfy
  1. 1.
    $\boldsymbol{Q}\in\mathcal{S}^{\text{doub}}$.
  2. 2.
    If the input $\boldsymbol{P}\in\mathcal{S}_{\varepsilon}$, we have $\boldsymbol{Q}\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$.
  3. 3.
    $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$.
  4. 4.
    $\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}\leq 2\left(\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\right)$.
Proof:
See Appendix A. ∎
Next, we introduce the following lemma that bounds the difference $\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)$.
Lemma 3
We have $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ and
$\lVert\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)\rVert_{1}\leq 2\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert_{1}.$
Proof:
The fact that $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ follows directly from Lemma 2-2.
For the other part, we only consider the case when $t$ is even. The case when $t$ is odd follows similarly.
Claim 1: When $t$ is even, we have
$\lVert\boldsymbol{P}(t)-\boldsymbol{\tilde{P}}(t)\rVert_{1}\leq 2\lVert\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}.$
This again follows by direct application of Lemma 2-4. (Notice that $\lVert\boldsymbol{\tilde{P}}(t)\boldsymbol{1}-\boldsymbol{1}\rVert_{1}=0$, since $\boldsymbol{\tilde{P}}(t)\in\mathcal{S}^{\text{row}}$ when $t$ is even)
Claim 2: When $t$ is even, we have
$\lVert\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\leq\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert_{1}.$
Notice that
$\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert_{1}$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}|\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)|\geq_{(a)}\sum_{j=1}^{s}\left\lvert\sum_{i=1}^{s}\tilde{P}_{i,j}(t+1)-\sum_{i=1}^{s}\tilde{P}_{i,j}(t)\right\rvert$
$=\lVert\boldsymbol{\tilde{P}}(t+1)^{\top}\boldsymbol{1}-\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}\rVert_{1}=\lVert\boldsymbol{\tilde{P}}(t)^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1},$
where (a) follows from the triangle inequality and the last equality follows since $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{S}^{\text{col}}$ when $t$ is even.
Combining the two claims, we are done. ∎
Now, we introduce some useful preliminary lemmas. First, we introduce the push-back lemma regarding minimizing strongly convex functions (See, for example [51]).
Lemma 4 (Push-back)
Consider $\mathcal{Q}\in\{\mathcal{S}^{\text{doub}}_{\varepsilon},\mathcal{S}^{\text{row}}_{\varepsilon},\mathcal{S}^{\text{col}}_{\varepsilon},\mathcal{S}\}$, and $\varepsilon\in[0,1/s]$. Let $g:\mathbb{R}_{+}^{s\times s}\to\mathbb{R}$ be a convex function. Fix $\alpha>0$, and $\boldsymbol{Y}\in\mathcal{Q}$. Let
$\boldsymbol{X}^{*}\in\arg\min_{\boldsymbol{X}\in\mathcal{Q}}\left[g(\boldsymbol{X})+\alpha D(\boldsymbol{X}\lVert\boldsymbol{Y})\right],$
where $D(\cdot\lVert\cdot)$ is the divergence defined in (15). Then,
$g(\boldsymbol{X}^{*})+\alpha D(\boldsymbol{X}^{*}\lVert\boldsymbol{Y})\leq g(\boldsymbol{Z})+\alpha D(\boldsymbol{Z}\lVert\boldsymbol{Y})-\alpha D(\boldsymbol{Z}\lVert\boldsymbol{X}^{*}),$
for all $\boldsymbol{Z}\in\mathcal{Q}$.
We now establish the following two lemmas.
Lemma 5
We have for $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}$,
$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert^{2},$ (20)
where $D(\boldsymbol{X}\lVert\boldsymbol{Y})$ is the divergence defined in (15).
Proof:
Notice that
$D(\boldsymbol{X}\lVert\boldsymbol{Y})\geq\sum_{i=1}^{s}\frac{1}{2}\left(\sum_{j=1}^{s}|X_{i,j}-Y_{i,j}|\right)^{2}\geq\frac{1}{2s}\lVert\boldsymbol{X}-\boldsymbol{Y}\rVert_{1}^{2}$
where the first inequality follows from the Pinsker’s inequality and the last inequality follows from the fact that for $a_{1},a_{2},\dots,a_{s}\in\mathbb{R}$, $\sum_{i=1}^{s}a_{i}^{2}\geq\left(\sum_{i=1}^{s}a_{i}\right)^{2}/s$. ∎
Next, we state a result on the properties of divergence D.
Lemma 6
We have $D(\boldsymbol{X}\lVert\boldsymbol{Y})\leq s\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{X}\in\mathcal{S}$, where $\boldsymbol{Y}\in\mathcal{S}_{\varepsilon}$.
Proof:
Notice that
$D(\boldsymbol{X}\lVert\boldsymbol{Y})$ $=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right)\leq_{(a)}\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{1}{\varepsilon}\right)=s\log\left(\frac{1}{\varepsilon}\right),$
where (a) follows since $X_{i,j}\leq 1$, $Y_{i,j}\geq\varepsilon$, and log is a non-decreasing function. ∎
Before moving on to the main theorem, we establish a deterministic bound on the queue size $\lVert\boldsymbol{Q}(t)\rVert$ for our Algorithm (Algorithm 1).
Lemma 7
We have for all $t\in\mathbb{N}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$, where $B$ is the bound on the subgradients of $\phi$ defined in assumption A1 and $V$ is the utility parameter of Algorithm 1.
Proof:
We first prove the following claim.
Claim: If $Q_{k}(t)>BV$, for $k\in[n]$, then $\gamma_{k}(t+1)=0$.
Proof:
Assume the contrary $Q_{k}(t)>BV$, and $\gamma_{k}(t+1)>0$.
Consider $\boldsymbol{\tilde{\gamma}}$ such that
$\tilde{\gamma}_{i}=\begin{cases}0&\text{ if }i=k\\ \gamma_{i}(t+1)&\text{ otherwise.}\end{cases}$
Since, $Q_{k}(t)>BV$, we have
$[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (21)
due to the bounded subgradient property (Assumption A1). Hence, notice that
$\sum_{i=1}^{n}$ $[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})=\gamma_{k}(t+1)[-V\phi^{{}^{\prime}}_{k}(\boldsymbol{\tilde{\gamma}})+Q_{k}(t)]>0,$ (22)
where the last inequality follows due to $\gamma_{k}(t+1)>0$ from assumption and (21).
The subgradient inequality for the convex function $f:[0,1]^{n}\to\mathbb{R}$ given by, $f(\boldsymbol{\gamma})=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}$ gives,
$\sum_{i=1}^{n}[-V\phi^{{}^{\prime}}_{i}(\boldsymbol{\tilde{\gamma}})+Q_{i}(t)](\gamma_{i}(t+1)-\tilde{\gamma}_{i})\leq f(\boldsymbol{\gamma}(t+1))-f(\boldsymbol{\tilde{\gamma}})\leq 0,$
where the last inequality follows from the optimality of $\boldsymbol{\gamma}(t+1)$ for (P3). This is a contradiction with (22). ∎
Now, we move on to the proof of Lemma 7. Fix $i\in[n]$. We use induction to prove that $Q_{i}(t)\leq BV+1$ for all $t\in\mathbb{N}$. Notice that the result trivially follows for $t=1$. Assume $Q_{i}(t)\leq BV+1$. We establish $Q_{i}(t+1)\leq BV+1$. We consider two cases.
Case 1: $Q_{i}(t)>BV$: Notice that from the previous claim, we should have $\gamma_{i}(t+1)=0$. Hence, we have from the queuing equation (16),
$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)\leq BV+1,$
where (a) follows since $\gamma_{i}(t+1)=0$, and $X_{i}(t)\geq 0$, and the last inequality follows from induction hypothesis.
Case 2: $Q_{i}(t)\leq BV$: We have from the queuing equation (16),
$Q_{i}(t+1)$ $\leq Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\leq_{(a)}Q_{i}(t)+1\leq BV+1,$
where (a) follows since $\gamma_{i}(t+1)\leq 1$, and $X_{i}(t)\geq 0$, and the last inequality follows from the description of the case.
Hence, we are done. ∎
Define the drift $\Delta(t)$ as
$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
and the history $\mathcal{H}(t)$ as the sigma algebra generated by
[…]
Notice that $\boldsymbol{\tilde{P}}(\tau),\boldsymbol{P}(\tau),\boldsymbol{\gamma}(\tau),\boldsymbol{Q}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable.
Now, we have the following bound on $\Delta(t)$.
Lemma 8
We have for all $t\in\mathbb{N}$,
$\Delta(t)$ $\leq n+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
Proof:
Notice that from the queuing equation (16), we have for all $i\in[n]$,
$Q^{2}_{i}(t+1)$ $\leq\left(Q_{i}(t)+\gamma_{i}(t+1)-X_{i}(t)\right)^{2}$
$\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t+1)+X^{2}_{i}(t)+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right]$
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right],$
where for the last inequality, we have used $X_{i}(t)\in\{0,1\}$. Summing the above for $i\in[n]$, we have
$\lVert\boldsymbol{Q}(t+1)\rVert^{2}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t+1)-X_{i}(t)\right].$
Taking the expectations conditioned $\mathcal{H}(t)$:
$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\mathbb{E}\{\gamma_{i}(t+1)|\mathcal{H}(t)\}-\sum_{j=1}^{m}q_{i,j}P_{i,j}(t)\right].$
Taking expectations, and performing simple algebraic manipulations, we have the result. ∎
Lemma 9
We have
$-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}$ $\leq\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}$
$\ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$
where $\hat{S}_{i,j}(t)$ is defined in line 6 of Algorithm 1.
Proof:
Notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}+\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}.$ (24)
Now, we handle each of the above four terms separately. Notice that
$\sum_{i=1}^{n}$ $\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t+1)-\tilde{P}_{i,j}(t)]\right\}$
$\leq_{(a)}\frac{3\eta s}{2}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{6s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}\right\}$
$\leq\frac{3\eta s}{2}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{3\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\},$
where for (a) we use $\frac{3\eta s}{2}\lVert\boldsymbol{a}\rVert^{2}+\frac{1}{6\eta s}\lVert\boldsymbol{b}\rVert^{2}\geq\sum_{i=1}^{k}a_{i}b_{i}$ for $k$-dimensional vectors $\boldsymbol{a,b}$, and the last inequality follows from Lemma 5. Next, Notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}$$\left\{Q_{i}(t)\hat{S}_{i,j}(t)[\tilde{P}_{i,j}(t)-P_{i,j}(t)]\right\}$
$\leq 3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{12s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t)-\boldsymbol{P}(t)\rVert^{2}_{1}\right\}$
$\leq_{(a)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{3s\eta}\mathbb{E}\left\{\lVert\boldsymbol{\tilde{P}}(t+1)-\boldsymbol{\tilde{P}}(t)\rVert^{2}_{1}\right\}$
$\leq_{(b)}3\eta s\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{2}{3\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\},$
where (a) follows from Lemma 3, and (b) follows from Lemma 5. Next, notice that
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}(t)\right\}$ $=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}P_{i,j}(t)Y_{i,j}(t)\right\}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\},$
Now, using the above in (9) we have
$\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\}$
$\leq\frac{9\eta s}{2}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}$
$\ \ \ \ +\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q_{i}(t)P_{i,j}(t)\right\}.$ (25)
But notice that
$\mathbb{E}$$\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}Q^{2}_{i}(t)\hat{S}^{2}_{i,j}(t)\right\}=\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i,j}(t)}{P^{2}_{i,j}(t)}Y_{i,j}(t)\right\}$
$\leq_{(a)}\frac{s}{\varepsilon}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i,j}(t)}{P_{i,j}(t)}Y_{i,j}(t)\right\}=\frac{s}{\varepsilon}\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}\leq_{(b)}\frac{s}{\varepsilon}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}$
$\leq\frac{nms}{\varepsilon}(BV+1)^{2},$
where (a) follows since $P_{i,j}(t)\geq\varepsilon/s$ from Lemma 3, (b) follows since $q_{i,j}\leq 1$, and the last inequality follows due to Lemma 7. Combining with (9), we are done. ∎
Now, we introduce the following lemma that will be useful in deriving the final bounds.
Lemma 10
We have for any $T,T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
$VT\phi(\boldsymbol{\gamma})$ $-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)$
$\ \ \ \ +\frac{n(BV+1)^{2}}{2}.$
Proof:
Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 8, we have
$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}\mathbb{E}\{Q_{i}(t)P_{i,j}(t)\}$
$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}$
$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\right\}-\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\right\},$ (26)
where (a) follows by Lemma 9. Now notice that from the optimality of $\boldsymbol{\tilde{P}}(t+1),\boldsymbol{\gamma}(t+1)$ in (8) (see Algorithm 1) with Lemma 4, we have for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\varepsilon}$,
$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)$
$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{m}\hat{S}_{i,j}(t)P_{i,j}\right]+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))-\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1)).$
Taking expectations of the above, we have
$-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{\tilde{P}}(t+1)\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ \ \ \ -\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)\tilde{P}_{i,j}(t+1)\}$
$\ \ \ \ \leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\gamma_{i}\}-\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\{Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
$\ \ \ \ =-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\sum_{j=1}^{m}q_{i,j}P_{i,j}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}$
Substituting the above in (10), we have
$\Delta$ $(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n+\frac{9\eta nms^{2}}{2\varepsilon}(BV+1)^{2}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t+1))\}.$
Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T+T_{0}-1\}$, we have
$\frac{1}{2}$ $\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(T_{0}))\right\}-\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(T+T_{0}))\right\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left[\gamma_{i,j}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ \ +\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right),$ (27)
where for the last inequality, we have used Lemma 6. Rearranging the above, and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ from Lemma 7, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Next, we have the following lemma that combines iterates $\boldsymbol{\gamma}(T_{0}+1),\boldsymbol{\gamma}(T_{0}+2)\dots,\boldsymbol{\gamma}(T+T_{0})$, and $\boldsymbol{X}(T_{0}),\boldsymbol{X}(T_{0}+1)$ $\dots,\boldsymbol{X}(T+T_{0}-1)$.
Lemma 11
We have for any $T,T_{0}\in\mathbb{N}$,
$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)\leq\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T}.$
Proof:
Notice that
$\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}+1}^{T+T_{0}}\boldsymbol{\gamma}(t)\right)\right)$
$=\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right)\right)$
$\leq_{(a)}\phi\left(\frac{1}{T}\Bigg{(}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}[\boldsymbol{Q}(t+1)-\boldsymbol{Q}(t)]\right)\Bigg{)}$
$\leq_{(b)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)+\frac{\boldsymbol{Q}(T+T_{0})}{T}\right)$
$\leq_{(c)}\phi\left(\frac{1}{T}\left(\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right)+\frac{\sqrt{n}B\lVert\boldsymbol{Q}(T+T_{0})\rVert}{T},$
where (a) follows from the entrywise nondecreasing property of $\phi$, and the queuing equation (16), for (b) we use $\boldsymbol{Q}(T_{0})\geq 0$ and the entrywise nondecreasing property of $\phi$, and (c) follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.  [^2] Assumption A1 ensures that $\phi$ can be extended to an entrywise nondecreasing function that is $\sqrt{nB}$-Lipschitz continuous over the domain $[0,\infty)^{n}$.
Now, we are ready to establish the performance bound of the algorithm.
Theorem 1
For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/s)$, $\eta,V>0$, Algorithm 1 yields
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)$ $\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}$
$\ \ \ \ +\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}.$
In particular
  1. 1.
    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
  2. 2.
    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
Proof:
Substituting $\boldsymbol{P}^{*}(1-\varepsilon s)+\varepsilon\boldsymbol{1},\boldsymbol{\gamma}^{*}$ in Lemma 10, where $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of (P2) defined in Lemma 1 and $\boldsymbol{1}$ is the all 1 matrix, we have
$VT\phi(\boldsymbol{\gamma}^{*})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\left(\gamma^{*}_{i}-\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}\right)\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ \ +\varepsilon\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}\sum_{j=1}^{m}\left(q_{i,j}sP^{*}_{i,j}-q_{i,j}\right)\mathbb{E}\{Q_{i}(t)\}+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
$\leq_{(a)}nT+\frac{9\eta nms^{2}T}{2\varepsilon}(BV+1)^{2}+\varepsilon nmsT(BV+1)+\frac{s}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2},$ (28)
where (a) follows since $\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\geq\gamma_{i}^{*}$ for all $i\in[n]$ from the feasibility of $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ for (P2), and Lemma 7. Now, we divide both sides of the above inequality by $VT$ and use the Jensen’s inequality to obtain
$\phi($ $\boldsymbol{\gamma}^{*})-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
$\leq_{(a)}\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}.$
Combining with Lemma 11, we have
$\phi($ $\boldsymbol{\gamma}^{*})-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq_{(a)}\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
$\leq_{(b)}\frac{n}{V}+\frac{9\eta nms^{2}}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{nB(BV+1)}{T},$
where (a) follows due to Lemma 11, and (b) follows due to Lemma 7. Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{\text{opt}}$. Using Jensen’s inequality, we are done. ∎
IV Discussion
IV-A Adaptiveness
A careful inspection of the proof of Theorem 1 shows that the result holds even if the success probabilities $q_{i,j}$ changed before time $T_{0}$, as long as they remained constant during $[T_{0}:T_{0}+T-1]$. Hence, the adaptiveness is satisfied.
IV-B Enforcing user fairness
Consider the case when $n\geq m$ (the number of users is at least the number of channels). Assume we fix a $\theta\in(0,1/s]$ such that each user is required to transmit at least $\theta$ fraction of the time on average on each channel and no user can transmit more than $(1-(n-m)\theta)$ fraction of the time. This enables faster adaptation to the new optimality point.
To see this, notice that for (P1), we require the additional constraints of
$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y_{i,j}(t)\}\geq\theta.$
for all $i\in[n]$ and $j\in[m]$ and
$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}\{Y_{i,j}(t)\}\leq 1-(n-m)\theta,$
for all $i\in[n]$. This will simply transform the constraint $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$ of (P2) to $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\theta}$. Setting $\varepsilon=\theta$ in Algorithm 1, this will allow us to use $\boldsymbol{P}^{*}$ directly in (1) in Theorem 1, instead of $(1-\varepsilon s)\boldsymbol{P}^{*}+\varepsilon\boldsymbol{1}$, which gives
$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\theta V}(BV+1)^{2}+\frac{\theta nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\theta}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{nB(BV+1)}{T}.$
Now, since we assume $\theta$ to be a constant, using $\eta=\Theta(1/T)$, and $V=\Theta(\sqrt{T})$, we have
$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}=O\left(\frac{1}{\sqrt{T}}\right)$
IV-C Case $m=1$
This case can be approached using the work on bandits with vector rewards and concave utility functions [26]. However, the algorithms proposed are not adaptive. It turns out that the algorithm for the general multi-channel case can be simplified for the single-channel case preserving the adaptiveness. In addition, the simplified algorithm has faster convergence. The idea of the algorithm is to find a distribution $\boldsymbol{p}(t)$ over the users in the $(t-1)$-th iteration, after which we sample the user to be assigned on the $t$-th iteration from $\boldsymbol{p}(t)$. Similar to the multi-channel case, we use auxiliary variables $\boldsymbol{\gamma}(t)\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(t)\in\mathbb{R}^{n}_{+}$ for all $t\in\mathbb{N}$. For notational convenience, for each $t\in\mathbb{N}$ and $i\in[n]$, we will use $q_{i},S_{i}(t)$ and $Y_{i}(t)$ instead of $q_{i,1},S_{i,1}(t)$ and $Y_{i,1}(t)$, respectively. The algorithm uses three parameters $\varepsilon\in(0,1/n)$, $V>0$, and $\eta>0$. See Algorithm 2 for details of the implementation.
1 Initialize $\boldsymbol{p}(1)\in\Delta_{n,\varepsilon}$, $\boldsymbol{\gamma}(1)\in[0,1]^{n}$, and the virtual queues $\boldsymbol{Q}\in[0,BV+1]^{n}$.
2 for each iteration $t\in[1:T]$ do
3       Sample $a_{t}\sim\boldsymbol{p}(t)$ and set $\boldsymbol{Y}(t)=\boldsymbol{e}_{a_{t}}$ where $\boldsymbol{e}_{i}$ is the unit vector with $i$-th entry being 1.
4       Receive feedback $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$.
5       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using, $\hat{S}_{i}(t)=\frac{S_{i}(t)Y_{i}(t)}{p_{i}(t)}$ for all $i\in[1:n]$.
6       Find, $\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1)$ by solving the problem,
$(\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1))=\arg\min_{\boldsymbol{p}\in\Delta_{n,\varepsilon},\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})$
$+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\Big{]},$ (29)
where $D_{\text{KL}}$ is the KL-divergence.
7       Update the virtual queues,
$\boldsymbol{Q}(t+1)=[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)]_{+},$ (30)
where $\boldsymbol{X}(t)=\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
8
Algorithm 2 Single Channel Adpative MAC
Most of the results required for the error analysis are borrowed from the multi-channel case. We only prove the results that are unique to this case. We first focus on solving the intermediate problem (6). Notice that the problem can be separated into the problem of finding optimal $\boldsymbol{\gamma}(t+1)$ and $\boldsymbol{p}(t+1)$.
Finding $\boldsymbol{\gamma}(t+1)$: This is the same problem as in the multi-channel case.
Finding $\boldsymbol{p}(t+1)$: This is (P3) with $\boldsymbol{x}=\eta\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)$, and $\boldsymbol{y}=\boldsymbol{p}(t)$.
Similar to Lemma 5, we have Pinsker’s inequality for KL divergence.
Lemma 12 (Pinsker’s inequality)
For $\boldsymbol{x},\boldsymbol{y}\in\Delta_{n}$, we have that,
$D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert_{1}^{2}\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert^{2},$
We also have the following lemma similar to Lemma 6.
Lemma 13
We have $D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\leq\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{x}\in\Delta_{n}$ where $\boldsymbol{y}\in\Delta_{n,\varepsilon}$.
Notice that since the queue updates and the problem to solve $\boldsymbol{\gamma}(t+1)$ are the same as in the multi-channel case, we have the same queue bound.
Lemma 14
We have that for all $t\in\{1,2,,\dots\}$ and $i\in\{1,2\dots,n\}$, $Q_{i}(t)\leq BV+1$.
Define the drift $\Delta(t)$ as,
$\Delta(t)=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}.$ (31)
We have the following lemma.
Lemma 15
We have that for all $t\in\{1,2,\dots\}$,
$\Delta(t)\leq n+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$
Proof:
Proof follows repeating the same arguments as Lemma 8. ∎
Lemma 16
We have that,
-- $\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
$\ \ \ \ \leq\frac{\eta n}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$
where $\hat{S}_{i}(t)$ is defined in line 5 of Algorithm 2.
Proof:
Notice that,
$\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\}$
$=\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)[p_{i}(t+1)-p_{i}(t)]\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
$\leq_{(a)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{2\eta}\mathbb{E}\left\{\lVert\boldsymbol{p}(t+1)-\boldsymbol{p}(t)\rVert^{2}\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
$\leq_{(b)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$ (32)
where (a) follows from $\frac{1}{2\eta}\lVert\boldsymbol{a}\rVert^{2}+\frac{\eta}{2}\lVert\boldsymbol{b}\rVert^{2}\geq\sum_{i=1}^{n}a_{i}b_{i}$ for $n$-dimensional vectors $\boldsymbol{a,b}$, (b) follows from Lemma 12. Define the history $\mathcal{H}(t)$ similar to multi-channel case (see, (23)). Notice that $\boldsymbol{p}(\tau),\boldsymbol{Q}(\tau),\boldsymbol{\gamma}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable. Now, we handle the terms of (16) separately. Notice that,
$\mathbb{E}$$\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}^{2}(t)\right\}$
$=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}\hat{S}_{a_{t}}(t)\right\}\leq_{(a)}\frac{1}{\varepsilon}\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}(t)\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\right\}$
$=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=_{(b)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}\mathbb{E}\left\{S_{i}(t)Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}$
$=_{(c)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}q_{i}\mathbb{E}\left\{Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}\leq_{(d)}\frac{1}{\varepsilon}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
$\leq\frac{n}{\varepsilon}(BV+1)^{2}$
where $a_{t}$ is defined in line 3 of Algorithm 2, (a) follows since $S_{a_{t}}(t)\leq 1$ and $p_{a_{t}}(t)\geq\varepsilon$, (b) follows since $\boldsymbol{Q}(t)$ and $\boldsymbol{p}(t)$ are $\mathcal{H}(t)$-measurable, (c) follows since $\boldsymbol{S}(t)$ is independent of $\boldsymbol{Y}(t)$ and $\mathcal{H}(t)$, (d) follows since $q_{i}\leq 1$, and the last inequality follows due to Lemma 14. Also,
$\sum_{i=1}^{n}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}=\mathbb{E}\left\{Q_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}p_{a_{t}}(t)\right\}=\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
Combining everything and substituting in (16), we are done. ∎
Now, we introduce the following lemma.
Lemma 17
We have that for any $T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
$VT\phi(\boldsymbol{\gamma})-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
Proof:
Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 15, we have that,
$\Delta(t)$ $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$
$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{\eta n}{2\varepsilon}(BV+1)^{2}$
$\ \ \ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$ (33)
where (a) follows by Lemma 16. Now notice that combining the decision (6) with Lemma 4, we have that for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))-\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)$
$\ \ \ \ \leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))-\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1)).$
Taking expectations of the above, we have that,
$-V$ $\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ \ -\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}p_{i}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}.$
Substituting the above in (17), we have that,
$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n+\frac{\eta n}{2\varepsilon}(BV+1)^{2}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, we have that,
$\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\}-\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\right\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right),$
where for the last inequality, we have used Lemma 13. Rearranging the above and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Now, we are ready to establish the performance bound of the algorithm.
Theorem 2
For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/n)$, $\eta,V>0$, we have that,
$\phi^{\text{opt}}$ $-\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\boldsymbol{X}(t)\}\right)$
$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
In particular
  1. 1.
    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
  2. 2.
    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
Proof:
Consider the problem
$\underset{\boldsymbol{p},\boldsymbol{\gamma}}{\text{max }}\phi\left(\boldsymbol{\gamma}\right)$ (34)
[…]
$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (36)
$\ \ \ \ \ q_{i}p_{i}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (37)
Using an argument similar to Lemma 1, it can be shown that the optimal objective value of the above problem is $\phi^{\text{opt}}$. Let $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ denote the solution of the above problem. Substituting $\boldsymbol{p}^{*}(1-\varepsilon n)+\varepsilon\boldsymbol{1},\boldsymbol{\gamma}^{*}$ in Lemma 17 we have that,
$VT\phi^{\text{opt}}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i}(1-\varepsilon n)-\varepsilon q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i})\mathbb{E}\{Q_{i}(t)\}+\varepsilon\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}nq_{i}p^{*}_{i}\mathbb{E}\{Q_{i}(t)\}$
$\ \ \ \ +\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
$\leq_{(a)}nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\varepsilon n^{2}T(BV+1)+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (38)
where (a) follows since $q_{i}p_{i}^{*}\geq\gamma_{i}^{*}$ (since $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ is feasible for problem (34)-(37)), and Lemma 18. Now, we use the Jensen’s inequality to obtain,
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
Notice that Lemma 11 holds since the queue updates are the same as in the multi-channel case. Combining above with Lemma 11, we have that,
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq_{(a)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
$\leq_{(b)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
where (a) follows due to Lemma 11, and (b) follows due to Lemma 18. Now, using Jensen’s inequality, we are done. ∎
Comparing the error bounds of Theorem 2 and Theorem 1 (with $m=1$), it can be seen that the bound of Theorem 2 is strictly better. In particular, the second term of the error bound has a $\Theta(n)$ dependence on $n$ in Theorem 2, whereas the corresponding value is $\Theta(n^{3})$ in Theorem 1.
V UCB Algorithm
In this section, we present a UCB-based algorithm for the link selection problem in MAC (Algorithm 3), after which we move on to the analysis. Recall that $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ is the user-channel assignment during time slot $t$. For $t\in[s]$, we set $\boldsymbol{Y}(t)$ arbitrarily such that each user-channel pair is explored exactly once during the interval $[s]$. For each $t>s$, we compute a permutation matrix $\boldsymbol{M}(t)$, an auxiliary vector $\boldsymbol{\gamma}(t)$, and a virtual queue $\boldsymbol{Q}(t)=[Q_{1}(t),\dots,Q_{n}(t)]$ similar to Algorithm 1. The assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ for the $t>s$ is obtained by discarding the last $s-m$ columns or the last $s-n$ rows of $\boldsymbol{M}(t)$ depending on whether $n$ or $m$ is larger. The algorithm uses parameters $\delta_{s},\delta_{s+1},\dots$ that satisfy $\delta_{t}>0$ for all $t\in\{s,s+1,\dots\}$. For each $i\in[n]$, $j\in[m]$, we define,
$n_{i,j}(t)=\sum_{\tau=1}^{t}Y_{i,j}(t),\ \forall t\in\mathbb{N}$
$\hat{S}_{i,j}(t)=\begin{cases}\frac{\sum_{\tau=1}^{t}Y_{i,j}(t)S_{i,j}(t)}{n_{i,j}(t)}&\text{ if }n_{i,j}(t)>0\\ 0&\text{ otherwise}\end{cases},\ \forall t\in\mathbb{N}$
$f_{i,j}(t)=\sqrt{\frac{\log\left(\frac{n_{i,j}(t)[n_{i,j}(t)+1]}{\delta_{t}}\right)}{2n_{i,j}(t)}},\ \forall t\in\{s,s+1,\dots\}$
$\text{UCB}_{i,j}(t)=\hat{S}_{i,j}(t)+f_{i,j}(t),\ \forall t\in\{s,s+1,\dots\}$ (39)
1 For $t\in[s]$, choose $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ arbitrarily such that each user-channel link is explored exactly once.
2
3Initialize $\boldsymbol{Q}(s+1)=0$.
4 for each time slot $t\in\{s+1,s+2,\dots,\}$ do
5       Find $\boldsymbol{\gamma}(t)$ and $\boldsymbol{M}(t)$ by solving
$\boldsymbol{\gamma}(t)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{N}Q_{i}(t)\gamma_{i},$ (40)
$\boldsymbol{M}(t)=\arg\max_{\boldsymbol{M}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\text{UCB}_{i,j}(t-1)Y_{i,j},$
where $\text{UCB}_{i,j}(t-1)$ is defined in (V).
6       Use the assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$, where $Y_{i,j}(t)=M_{i,j}(t)$ for $(i,j)\in[n]\times[m]$ and receive feedback $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
7       Update the queues
$Q_{i}(t+1)=[Q_{i}(t)+\gamma_{i}(t)-X_{i}(t)]_{+},$ (41)
$\forall\ i\in[n]$, where $X_{i}(t)=\sum_{j=1}^{m}Y_{i,j}(t)S_{i,j}(t)$.
8
Algorithm 3 UCB MAC
We first focus on solving the inner problem (40). Notice that the problem to solve to obtain $\boldsymbol{\gamma}(t)$ is the same as for Algorithm 1. The problem to solve to obtain $\boldsymbol{M}(t)$ has a classic max-weight structure (hence, $\boldsymbol{M}(t)$ is a permutation matrix). Hence, we can use the Hungarian algorithm [49] to obtain the solution.
Now, we move on to the analysis of the Algorithm 3. Since the queue updates and the problem to solve to obtain $\boldsymbol{\gamma}(t)$ are the same as in Algorithm 1, we have the same deterministic queue bound. We formally state this in the following lemma.
Lemma 18
We have for all $t\in\{s+1,s+2,\dots\}$ and $i\in[n]$, $Q_{i}(t)\leq BV+1$
Proof:
This follows repeating the argument of Lemma 7. ∎
Define the collection of good events $\Omega_{s+1},\Omega_{s+2},\dots$ as
[…]
where $\text{UCB}_{i,j}(t-1)$ and $f_{i,j}(t-1)$ are defined in (V). We have the following lemma.
Lemma 19
For each $t\in\{s+1,s+2,\dots,\}$, we have $\mathbb{P}\{\Omega_{t}\}\geq 1-2nm\delta_{t-1}$.
Proof:
The proof follows directly by applying the Hoeffding inequality with union bound. See [48] for more details. ∎
Define the drift $\Delta_{t}$
$\Delta_{t}=\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}-\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
and history $\mathcal{H}(t)$ as
[…]
for each $t\in\{s+1,s+2,\dots\}$. Notice that $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable.
Lemma 20
We have for each $t\in\{s+1,s+2,\dots\}$
$\Delta(t)\leq n+\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\right\}.$
Proof:
Proof is similar to the proof of Lemma 8 and is given in Appendix B. ∎
Fix a time horizon $T\in\{s+1,s+2,\dots\}$. Then we have the following result.
Lemma 21
Given a time horizon $T\in\{s+1,s+2,\dots\}$, $V>0$, and $\delta_{s},\delta_{s+1},\dots,\delta_{T-1}>0$, we have
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{\gamma}(t)\right)\right\}$ $\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}$
$\ \ \ \ +\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t},$
where $f_{i,j}(t)$ is defined in (V).
Proof:
We begin with the following two claims.
Claim 1: We have that,
$-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\leq V\phi^{\text{max}}+nBV+n$
Proof:
The proof immediately follows from Lemma 18 and the definition of $\phi^{\text{max}}$ in Assumption A1. ∎
Claim 2: We have that,
$\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\Bigg{|}\Omega_{t}\right\}$
$\leq-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
Proof:
From the definition of the good event $\Omega_{t}$ in (42), we have
$\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\geq\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}-2\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\geq\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)M_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\ \ \ \ -2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$ (44)
where the last inequality follows from Lemma 18. Also, notice that,
$\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}(t)\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)M_{i,j}(t)\Bigg{|}\Omega_{t}\right\}$
$\leq_{(a)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}\text{UCB}_{i,j}(t-1)P_{i,j}^{*}\Bigg{|}\Omega_{t}\right\}$
$\leq_{(b)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\Bigg{|}\Omega_{t}\right\}\leq-V\phi^{\text{opt}}$ (45)
where (a) follows from the optimality of $\boldsymbol{\gamma}(t),\boldsymbol{M}(t)$ for intermediate problem (40) in Algorithm 3 (recall $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of P2); (b) follows from the definition of the good event $\Omega_{t}$ in (42); the last inequality follows from (11). Adding the inequalities (21) and (21) and rearranging, we are done. ∎
Combining claim 1 and claim 2, and the law of total probability, we have that,
$\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}(t))+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right]\right\}$
$\leq(V\phi^{\text{max}}+nBV+n)\mathbb{P}\{\Omega_{t}^{c}\}-V\phi^{\text{opt}}\mathbb{P}\{\Omega_{t}\}$
$\ \ \ \ +2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\Bigg{|}\Omega_{t}\right\}\mathbb{P}\{\Omega_{t}\}$
$\leq_{(a)}(V\phi^{\text{max}}+V\phi^{\text{opt}}+nBV+n)\mathbb{P}\{\Omega_{t}^{c}\}-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$
$\leq_{(b)}2(2V\phi^{\text{max}}+nBV+n)nm\delta_{t-1}-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$ (46)
where (a) follows from the fact that for a nonnegative random variable $\boldsymbol{X}$ and an event $\mathcal{H}$ we have $\mathbb{E}\{X|\mathcal{H}\}\mathbb{P}\{\mathcal{H}\}\leq\mathbb{E}\{X\}$ and (b) follows since $\phi^{\text{opt}}\leq\phi^{\text{max}}$ and Lemma 19. Adding (21) to the result of Lemma 20 and rearranging, we have
$\Delta(t)-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t))\}$
$\leq n-V\phi^{\text{opt}}+2(BV+1)\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}+2(2V\phi^{\text{max}}+nBV+n)nm\delta_{t-1}$
Now, we sum the above for $t\in[s+1,T]$ to obtain
$\frac{1}{2}\mathbb{E}\{\lVert\boldsymbol{Q}(T+1)\rVert^{2}-\lVert\boldsymbol{Q}(s+1)\rVert^{2}\}-V\sum_{t=s+1}^{T}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t))\}$
$\leq n(T-s)-V(T-s)\phi^{\text{opt}}+2(BV+1)\sum_{t=s+1}^{T}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$
$\ \ \ \ +2(2V\phi^{\text{max}}+nBV+n)nm\sum_{t=s}^{T-1}\delta_{t}.$
Using $\lVert\boldsymbol{Q}(T+1)\rVert^{2}\geq 0$ and $\lVert\boldsymbol{Q}(s+1)\rVert^{2}=0$, we have
$V(T-s)\phi^{\text{opt}}-V\sum_{t=s+1}^{T}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t))\}$
$\leq n(T-s)+2(BV+1)\sum_{t=s+1}^{T}\mathbb{E}\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}f_{i,j}(t-1)Y_{i,j}(t)\right\}$
$\ \ \ \ +2(2V\phi^{\text{max}}+nBV+n)nm\sum_{t=s}^{T-1}\delta_{t}.$
Dividing both sides by $V(T-s)$, and using Jensen’s inequality, we are done. ∎
Now we have the following lemma that combines $\boldsymbol{\gamma}(s+1),\boldsymbol{\gamma}(s+2),\dots,\boldsymbol{\gamma}(T)$ with $\boldsymbol{X}(s+1),\boldsymbol{X}(s+2),\dots,\boldsymbol{X}(T)$.
Lemma 22
We have
$\phi\left(\frac{1}{T-s}\left(\sum_{t=s+1}^{T}\boldsymbol{\gamma}(t)\right)\right)\leq\phi\left(\frac{1}{T-s}\left(\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right)\right)+\frac{nB(BV+1)}{T-s}.$
Proof:
The proof uses the same argument as Lemma 11. ∎
Now, we have the following theorem that establishes the performance bound of Algorithm 3.
Theorem 3
Given a time horizon $T\in\{s+1,s+2,\dots\}$, $V>0$, and $\delta_{s},\delta_{s+1},\dots,\delta_{T-1}>0$, we have
$\phi^{\text{opt}}$ $-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
$\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t}$
$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T},$ (47)
where $f_{i,j}(t)$ is defined in (V). In particular using $\delta_{t}=1/t$ for all $t\geq s$, we have
  1. 1.
    For all $V>0$,
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
    $\leq\frac{n}{V}+\frac{2\sqrt{6}nm(BV+1)\sqrt{T\log(T)}}{V(T-s)}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm(\log(T)+1)}{V(T-s)}$
    $\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$
  2. 2.
    Assume $T\geq 2s$. Using $V=\Theta(\sqrt{T})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right),$
    and $\mathcal{O}$ hides the dependence on all parameters but $T$.
  3. 3.
    Fix $\epsilon>0$. Using $V=\Theta(1/\epsilon)$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}(\epsilon),$
    and $\mathcal{O}$ hides the dependence on all parameters but $\epsilon$.
Proof:
First, notice that
$\left\lvert\phi\left(\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right)-\phi\left(\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\rvert$
$\leq\sqrt{n}B\left\lVert\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{X}(t)-\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\rVert\sqrt{n}B\left\lVert\frac{1}{T}\sum_{t=1}^{s}\boldsymbol{X}(t)+\frac{s}{T(T-s)}\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right\rVert$
$\leq\frac{\sqrt{n}B}{T}\sum_{t=1}^{s}\left\lVert\boldsymbol{X}(t)\right\rVert+\frac{s}{T(T-s)}\sum_{t=s+1}^{T}\left\lVert\boldsymbol{X}(t)\right\rVert\leq\frac{\sqrt{n}sB}{T}+\frac{s}{T}=\frac{(\sqrt{n}B+1)s}{T},$ (48)
where the first inequality follows since $\phi$ is $\sqrt{n}B$-Lipschitz continuous from Assumption A1.
Combining Lemma 22 and Lemma 21, we have
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T-s}\sum_{t=s+1}^{T}\boldsymbol{X}(t)\right)\right\}$
$\leq\frac{n}{V}+\frac{2(BV+1)}{V(T-s)}\sum_{t=s+1}^{T}\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbb{E}\left\{Y_{i,j}(t)f_{i,j}(t-1)\right\}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\delta_{t}$
$\ \ \ \ +\frac{nB(BV+1)}{T-s}.$ (49)
Combining (3) and (3) and using the Jensen’s inequality, we have (3).
To prove part 1, first fix $i\in[n]$ and $j\in[m]$. When $\delta_{t-1}=1/(t-1)$ for all $t\in[s+1:T]$, we have
$f_{i,j}(t-1)$ $=\sqrt{\frac{\log\left(\frac{n_{i,j}(t-1)[n_{i,j}(t-1)+1]}{\delta_{t-1}}\right)}{2n_{i,j}(t-1)}}\leq\sqrt{\frac{3\log\left(T\right)}{2n_{i,j}(t-1)}}.$ (50)
Also, $n_{i,j}(t)-n_{i,j}(t-1)=1$ if and only if $Y_{i,j}(t)=1$. Hence, we have
$\mathbb{E}\left\{\sum_{t=s+1}^{T}\frac{Y_{i,j}(t)}{\sqrt{n_{i,j}(t-1)}}\right\}=\mathbb{E}\left\{\sum_{k=1}^{n_{i,j}(T-1)-1}\sqrt{\frac{1}{k}}\right\}\leq\mathbb{E}\left\{\sum_{k=1}^{T}\sqrt{\frac{1}{k}}\right\}\leq 2\sqrt{T}$ (51)
where for the last inequality we have used $\sum_{\tau=1}^{t}\frac{1}{\sqrt{\tau}}\leq 2\sqrt{t}$ for all $t\geq 1$. Combining (50) and (51), we have
$\sum_{t=s+1}^{T}\mathbb{E}\{Y_{i,j}(t)f_{i,j}(t-1)\}\leq\sqrt{6T\log(T)}$
Now substituting the above in (3), we have
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
$\leq\frac{n}{V}+\frac{2\sqrt{6}nm(BV+1)\sqrt{T\log(T)}}{V(T-s)}+\frac{2(2V\phi^{\text{max}}+nBV+n)nm}{V(T-s)}\sum_{t=s}^{T-1}\frac{1}{t}$
$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$ (52)
Part 1 follows from above after using the fact that $\sum_{\tau=1}^{t}\frac{1}{\tau}\leq\ln(t)+1$ for all $t\geq 1$.
For part 2, notice that since $T\geq 2s$, from part 1 we have
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)$
$\leq\frac{n}{V}+\frac{4\sqrt{6}nm(BV+1)\sqrt{\log(T)}}{V\sqrt{T}}+\frac{4(2V\phi^{\text{max}}+nBV+n)nm(\log(T)+1)}{VT}$
$\ \ \ \ +\frac{2nB(BV+1)}{T}+\frac{(\sqrt{n}B+1)s}{T},$
which is clearly $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$ when $V=\Theta(\sqrt{T})$.
For part 3, notice that from part 1,
$\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)\leq\frac{n}{V}.$
Using $V=\Theta(1/\epsilon)$, we are done. ∎
VI Discussion
VI-A Distributed Implementation
Consider a system where the $n$ users take decisions in the absence of a centralized controller. In each time slot, each user receives as feedback whether each channel was successfully accessed or not (this can be, for instance, via sensing each channel). Notice that $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$ are $\mathcal{H}(t)$-measurable, where $\mathcal{H}(t)$ is defined in (43). Hence, during time slot $t$, if each user has access to the full history $\mathcal{H}(t)$, they can independently implement Algorithm 3 to find $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$. From the feedback, each user can compute $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$ and hence the full history $\mathcal{H}(t+1)$. Hence, starting from $\mathcal{H}(s+1)$, each user can independently implement Algorithm 3. Notice that this is not possible in Algorithm 1, since the channel assignments are not $\mathcal{H}(t)$-measurable.
VII Simulations
For simulations, we consider four scenarios with $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+x_{i})$, $n=5,m=3$, and $T=10^{5}$. In each scenario, the values of $q_{i,j}$ change mid-way through the simulation (at $T=5\times 10^{4}$). In Scenario 1 (Figure 1-Left-Top), certain user-channel links are turned off at $T=5\times 10^{4}$. In Scenario 2 (Figure 1-Right-Top), certain user-channel links that were turned off before, turn on at $T=5\times 10^{4}$. In Scenarios 3 (Figure 1-Left-Bottom) and 4 (Figure 1-Right-Bottom), the values of $q_{i,j}$ for certain user-channel links change at $T=5\times 10^{4}$. In each scenario, the algorithms are not told about the change. For each figure, for each algorithm, we plot $\phi\left(\frac{1}{t}\sum_{\tau=1}^{t}\boldsymbol{X}(\tau)\right)$ for $t\leq T/2$ and $\phi\left(\frac{1}{t-T/2}\sum_{\tau=T/2+1}^{t}\boldsymbol{X}(\tau)\right)$ for $t>T/2$ vs $t$. We also plot the optimal objective value of (P2), for reference. Note that in the first half of the simulation, both algorithms converge to the optimal solution in all scenarios, although the convergence of Algorithm 3 is faster. However, to show the importance of Algorithm 1, we observe the second half of the simulation. Although Algorithm 3 achieves superior performance in the first half of the simulation, it does not perform well in the second half except in Scenario 2. In contrast Algorithm 1 achieves similar performance in both halves in all scenarios. This is due to the adaptiveness of Algorithm 1.
Figure 1: Top-Left: Scenario 1. Top-Right: Scenario 2. Bottom-Left: Scenario 3. Bottom-Right: Scenario 4.
VIII Conclusions
This paper focused on the problem of designing algorithms for automatic link selection in multi-channel multiple access with link failures. In particular, we solved a network utility maximization problem with matching constraints and bandit feedback on link failures. We considered two algorithms, where the first algorithm has slower convergence and is adaptive, and the second algorithm has faster convergence and is not adaptive.
Appendix A Proof of Lemma 2
Before proving this, notice that due to the normalization steps 1 and 2 of the ROUND function, we have $\boldsymbol{1}\geq\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}$, and $\boldsymbol{1}\geq(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}$, where the inequalities are taken entrywise. Hence, we have $\boldsymbol{Q}\geq\boldsymbol{P}^{{}^{\prime\prime}}\geq 0$. Now, we prove each part separately.
1) First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$. Then $\boldsymbol{P}^{{}^{\prime\prime}}\in\mathcal{S}^{\text{row}}$ and $\boldsymbol{Q}=\boldsymbol{P}^{{}^{\prime\prime}}$ (by the definition of $\boldsymbol{Q}$ in (19)). Also, due to the normalization of the columns in step 2, each column of $\boldsymbol{P}^{{}^{\prime\prime}}$ has a sum of at most 1. However, notice that since $\boldsymbol{P}^{{}^{\prime\prime}}\in\mathcal{S}^{\text{row}}$, the sum of its entries is $n$. Hence, the sum of each column must be exactly 1. Hence, $\boldsymbol{Q}=\boldsymbol{P}^{{}^{\prime\prime}}\in\mathcal{S}^{\text{doub}}$ as desired. Next, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}>0$. Then we have
$\boldsymbol{Q}\boldsymbol{1}$ $=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\boldsymbol{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=_{(a)}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{(\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}})\boldsymbol{1}}{\boldsymbol{1}^{\top}(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})}$
$=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}{\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}$
$=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{s\boldsymbol{1}-s\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}{\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}$
$=\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\frac{(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(s-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})}{s-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}=\boldsymbol{1},$
where in (a) we have used $\boldsymbol{1}\geq\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}$ when simplifying the denominator. The claim $\boldsymbol{1}^{\top}\boldsymbol{Q}=\boldsymbol{1}^{\top}$ follows repeating the same argument.
2) We will prove the case $\boldsymbol{P}\in\mathcal{S}^{\text{row}}_{\varepsilon}$. The other case follows similarly. Since $\boldsymbol{P}\in\mathcal{S}^{\text{row}}_{\varepsilon}$, from (17), we have $P^{{}^{\prime}}_{i,j}=P_{i,j}$. Also, since $\boldsymbol{P}\in\mathcal{S}^{\text{row}}_{\varepsilon}$ we have $P_{i,j}\leq 1$ for all $i,j\in\{1,\dots,s\}$. Hence $P^{{}^{\prime}}_{i,j}\leq 1$ for all $i,j\in\{1,\dots,s\}$. Hence, from (18), we have $P^{{}^{\prime\prime}}_{i,j}\geq P{{}^{\prime}}_{i,j}/s=P_{i,j}/s\geq\varepsilon/s$. Since we already know $\boldsymbol{Q}\geq\boldsymbol{P}^{{}^{\prime\prime}}$, we have $Q_{i,j}\geq\varepsilon/s$ for all $i,j$. Part 1 shows $Q\in\mathcal{S}^{\text{doub}}$, so $Q\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$. So we are done.
3) First, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}=0$. Then we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0$. Also, due to the scaling, we have $s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}\geq 0$. Hence, we have $\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}=0\leq s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ and we are done. Next, suppose $\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}>0$. Notice that
$\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$ $=\frac{\lVert(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=\frac{\lVert\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\lVert_{1}}{\lVert\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\rVert_{1}}$
$=_{(a)}\frac{\boldsymbol{1}^{\top}\left(\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}-\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}+\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\right)\boldsymbol{1}}{\boldsymbol{1}^{\top}\left(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\right)}$
$=\frac{\boldsymbol{1}^{\top}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}+\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}{\boldsymbol{1}^{\top}\boldsymbol{1}-\boldsymbol{1}^{\top}\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}}$
$=\frac{s^{2}-2s\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}^{2}}{s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}}=s-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1},$
where (a) follows since $\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1}\geq 0$, and $\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1}\geq 0$, as a result of which $(\boldsymbol{1}-\boldsymbol{P}^{{}^{\prime\prime}}\boldsymbol{1})(\boldsymbol{1}-(\boldsymbol{P}^{{}^{\prime\prime}})^{\top}\boldsymbol{1})^{\top}\geq 0$.
4) Denote by $\mathcal{U}=\left\{i\in[s]:\sum_{j=1}^{s}P_{i,j}>1\right\}$ and $\mathcal{U}^{c}=[s]\setminus\mathcal{U}$. Notice that
$\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}+\lVert\boldsymbol{P}\rVert_{1}-s$
$=\sum_{i=1}^{s}\left\lvert\sum_{j=1}^{s}P_{i,j}-1\right\rvert+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
$=\sum_{i\in\mathcal{U}}\left(\sum_{j=1}^{s}P_{i,j}-1\right)+\sum_{i\in\mathcal{U}^{c}}\left(1-\sum_{j=1}^{s}P_{i,j}\right)+\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}-s$
$=2\sum_{i\in\mathcal{U}}\sum_{j=1}^{s}P_{i,j}-|\mathcal{U}|+|\mathcal{U}^{c}|-s$
$=2\left(\sum_{i\in\mathcal{U}}\sum_{j=1}^{s}P_{i,j}-|\mathcal{U}|\right)=2\sum_{i\in\mathcal{U}}\left(\sum_{j=1}^{s}P_{i,j}-1\right)$
$=_{(a)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime}}\rVert_{1})=2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2(\lVert\boldsymbol{P}^{{}^{\prime}}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})$
$=_{(b)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left[\sum_{i=1}^{s}P^{{}^{\prime}}_{i,j}-1\right]_{+}\right)$
$\geq_{(c)}2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left[\sum_{i=1}^{s}P_{i,j}-1\right]_{+}\right)$
$\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\left(\sum_{j=1}^{s}\left\lvert\sum_{i=1}^{s}P_{i,j}-1\right\rvert\right)$
$\geq 2(\rVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1})-2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}$ (53)
where (a) and (b) follow from the definitions of $\boldsymbol{P}^{{}^{\prime}},\boldsymbol{P}^{{}^{\prime\prime}}$ in (17) and (18), respectively; (c) follows since $P_{i,j}\geq P^{{}^{\prime}}_{i,j}$ due to scaling. Rearranging the above inequality, we have
$\lVert\boldsymbol{P}\rVert_{1}-2\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+s\leq\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}.$ (54)
Now, to complete the proof notice that
$\lVert\boldsymbol{P}-\boldsymbol{Q}\rVert_{1}$ $\leq\lVert\boldsymbol{P}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
$=_{(a)}\lVert\boldsymbol{P}\rVert_{1}-\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+\lVert\boldsymbol{Q}-\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}$
$\leq_{(b)}\lVert\boldsymbol{P}\rVert_{1}-2\lVert\boldsymbol{P}^{{}^{\prime\prime}}\rVert_{1}+s\leq_{(c)}\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+2\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}$
$\leq 2\left(\lVert\boldsymbol{P}\boldsymbol{1}-\boldsymbol{1}\lVert_{1}+\lVert\boldsymbol{P}^{\top}\boldsymbol{1}-\boldsymbol{1}\rVert_{1}\right),$ (55)
where (a) follows since $\boldsymbol{P}^{{}^{\prime\prime}}\leq\boldsymbol{P}$ due to the scaling, (b) follows due to part 3, and (c) follows from (54).
Appendix B Proof of Lemma 20
Notice that from the queuing equation (41), we have
$Q^{2}_{i}(t+1)$ $\leq(Q_{i}(t)+\gamma_{i}(t)-X_{i}(t))^{2}\leq Q^{2}_{i}(t)+\gamma_{i}^{2}(t)+X^{2}_{i}(t)+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
for all $i\in[n]$. Summing the above for $i\in[n]$, we have
$\lVert\boldsymbol{Q}(t+1)\rVert^{2}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)].$
Taking the expectations conditioned on the history $\mathcal{H}(t)$ defined in (43), we have
$\mathbb{E}\{\lVert\boldsymbol{Q}(t+1)\rVert^{2}|\mathcal{H}(t)\}$ $\leq\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)[\mathbb{E}\{\gamma_{i}(t)|\mathcal{H}(t)\}-\mathbb{E}\{X_{i}(t)|\mathcal{H}(t)\}]$
$=\lVert\boldsymbol{Q}(t)\rVert^{2}+2n+2\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}(t)-\sum_{j=1}^{m}q_{i,j}Y_{i,j}(t)\right],$
where the last equality follows since $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable. Taking expectations, we have the result.